OpenStudy (anonymous):

A ball is thrown from an initial height of 7 feet with an initial upward velocity of 21 ft/s . The ball's height h (in feet) after t seconds is given by the following.

OpenStudy (anonymous):

\[h=7+21t-16t ^{2}\]

OpenStudy (anonymous):

Find all values of t for which the ball's height is 13 ft.

OpenStudy (anonymous):

Hint. All of the information and numbers given to you in the first sentence are irrelevant to answering the question. Only the second sentence and the equation are relevant.

OpenStudy (anonymous):

am I substituting t for 13?

OpenStudy (anonymous):

13ft is a height. t represents time. no. try again.

OpenStudy (anonymous):

do I replace h for 13 so it'd be \[13=7+21t-16t ^{2}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok. so then do I subtract 13 from both sides?

OpenStudy (anonymous):

that's right.

OpenStudy (anonymous):

ok. thennnnn do I factor? how do i solve for t?

OpenStudy (anonymous):

oh i see. alright i'll do that

OpenStudy (anonymous):

actualy, i would not subtract 13. I'd move everything over to the left side to get 16t^2 - 21t+ 6 = 0

OpenStudy (anonymous):

ok. what's next?

OpenStudy (anonymous):

And that, I'm afraid, is one ugly quadratic equation! It doesn't even factor according to my calculator..(my calculator can factor quadratics)..so you'll have to use the quadratic formula

OpenStudy (anonymous):

ok ok. i can do that, just give me a sec

OpenStudy (anonymous):

you should get a ti-89 calculator i can just type "solve16t^2-21t+6=0,t)" and it spits out the answers. lol

OpenStudy (anonymous):

I got 0.8921823261 or 0.4203176739

OpenStudy (anonymous):

that's what i got, too. good luck!

OpenStudy (anonymous):

oh that's it? that's t?

OpenStudy (anonymous):

t = 0.892 seconds and t = 0.420 seconds

OpenStudy (anonymous):

yup, that's what I have. Thank you so much!